the length of the longest side of a triangle is double that of the next longest side and 19cm greater than the shortest side. The perimeter is 41cm. What is the shortest side
Answer by mananth(12270)

let second side be =x

longest side = 2x
shortest side = 2x-19
x+2x+2x-19=41
5x-19=41
5x=60
x=12
shortest side = 2x-19
24-19=5
5+12+24 =41